# Towards Dataset Copyright Evasion Attack against Personalized Text-to-Image Diffusion Models

### v1

@@ IV-D Trigger Identification
[…]
@@ IV-E Efficient Watermark Mitigation
By updating the cross-attention weights using this expression, we can effectively erase the model’s sensitivity to specific triggers without degrading its performance on normal inputs. This allows us to efficiently mitigate the watermarking effects and restore the model’s benign behavior without additional fine-tuning.
V Experiments
V-A Main Settings
Datasets and Models. We adopt three benchmark datasets to evaluate all dataset copyright evasion attacks, i.e., Pokemon [1], Ossaili [2], and Pranked03 [3] datasets. All experiments are conducted using Stable Diffusion v1.4, which serves as our default T2I diffusion model.
Settings for DOV. We conduct four backdoor-based dataset ownership verifications, including BadT2I-Local (BadT2I-L) [63], BadT2I-Global (BadT2I-G) [63], Rickrolling [49], and Villan Diffusion (VD) [9]. For the text trigger, BadT2I-L and BadT2I-G use the word “university” as the trigger. Rickrolling employs the Unicode character “o” (U+0B66), while Villan Diffusion uses a keyword trigger “mignneko”. For the owner-specified target image, BadT2I-L is a $128\times 128$ local patch placed at the top-left corner of generated images. BadT2I-G and Rickrolling use a $512\times 512$ global target image, i.e., a Hello Kitty image, while VD uses a $512\times 512$ global target image, i.e., a BabyKitty image. The watermarking rate is set as $\gamma=20\%$. We fully fine-tune the T2I diffusion models on these datasets by using Adam optimizer with a learning rate of $10^{-6}$ for $T_{total}=100$ epochs. The resolution of the generated image is $512\times 512$.
Settings for CEA. We compared our CEAT2I with four different dataset copyright evasion attacks, including ABL [29], NAD [30], TPD [8], and T2IShield [54]. ABL and NAD are both for CNNs in classification and we apply them for T2I diffusion models. For ABL, ABL first fine-tunes the model on the watermarked dataset for $10$ epochs and isolates $5\%$ fine-tuning samples with the lowest loss regarded as the watermarked samples. Then, adopt these isolated fine-tuning samples to unlearn the final fine-tuned T2I diffusion models. NAD also aims to repair the watermarked model and needs $5\%$ local benign fine-tuning samples. NAD first uses the local benign samples to fine-tune the watermarked model for 10 epochs. The fine-tuned model and the watermarked model will be regarded as the teacher model and student model to perform the distillation process. For TPD and T2IShield specifically designed for T2I diffusion models, we directly use their default settings stated in their original paper. Our CEAT2I performs watermarked sample detection at the early fine-tuning epoch $T_{e}=30$ and the detection thresholds are set to $\alpha_{1}=0.4$ and $\alpha_{2}=15$. The trigger identification is conducted using the second-to-last convolutional layer of the model.
Evaluation Metrics. To evaluate the effectiveness of our dataset ownership evasion attacks, we adopt two key metrics from [63]. Specifically, we train a ResNet18 classifier for each owner-specified target image to detect whether a generated image contains the backdoor-based watermark. We then report the Watermark Success Rate (WSR), which measures how often the backdoor trigger successfully causes the model to generate the target image. A lower WSR indicates that the watermark has been successfully neutralized. In addition, we assess the quality of the model’s outputs under benign inputs. To this end, we compute the CLIP similarity score, which is the cosine similarity between the CLIP embeddings of the generated images and their corresponding ground-truth images. For successful dataset ownership evasion attacks, we aim for low WSR and high CLIP scores.
TABLE I: The CLIP similarity between images (CLIP %) and watermark success rate (WSR %) of one baseline without attacks and five different CEA methods against four types of DOV methods across three datasets, including Pokemon, Ossaili, Pranked03 datasets. The best results among five CEA methods are highlighted in bold. In particular, we mark the failure cases (i.e., WSR $>10\%$) among five CEA methods in red.
@@ V-B Main Results
V-C Ablation Study
Discussions on Watermarked Sample Detection. We compare the effectiveness of watermarked sample detection across ABL [29], T2IShield [54], and our proposed CEAT2I. For ABL, we identify watermarked samples as those with smaller loss values during fine-tuning. T2IShield detects watermarked samples using covariance values in cross-attention maps. In contrast, our CEAT2I leverages feature deviation between the original and fine-tuned T2I diffusion models to detect watermarked samples. Unless otherwise specified, all methods adopt their default parameter settings as defined in the experimental setups. As shown in Table II, ABL achieves low detection accuracy, and T2IShield struggles to detect watermarked samples in BadT2I-L, where the owner-specified target is a small image patch. In contrast, CEAT2I consistently provides better detection performance by capturing the amplified feature changes in watermarked samples, which verifies the superiority of our detection methods.
@@ V-C Ablation Study
Discussions on Trigger Identification. We evaluate the accuracy of our trigger identification approach. Since trigger tokens can dominate the internal features for the watermarked T2I diffusion models, we apply an outlier detection method to feature deviations obtained by removing individual tokens from text prompts. Our method successfully identifies trigger tokens for four DOV methods across three datasets, achieving 100% accuracy when applied to previously detected watermarked samples.
TABLE III: The CLIP similarity between images (CLIP %) and watermark success rate (WSR %) of one baseline without attacks and five different CEA methods against four types of DOV methods on the Pokemon dataset under different watermarking rates. The best results among five CEA methods are highlighted in bold. In particular, we mark the failure cases (i.e., WSR $>10\%$) among five CEA methods in red.
@@
TABLE IV: The CLIP similarity between images (CLIP %) and watermark success rate (WSR %) of one baseline without attacks and five different CEA methods against four types of DOV methods on the Pokemon dataset under different trigger positions. The best results among five CEA methods are highlighted in bold. In particular, we mark the failure cases (i.e., WSR $>10\%$) among five CEA methods in red.
@@
V-D Resistance to Potential Adaptive Defense

### v2

@@ IV-D Trigger Identification
In the second stage, CEAT2I locates potential trigger tokens within the detected watermarked samples rather than discarding the samples directly. Since triggers induce the model to produce watermark-specific outputs that amplify feature divergence, CEAT2I identifies candidate trigger tokens by measuring feature deviations caused by the removal of individual words from the input prompts.
@@
$\hat{y}_{w}^{tr}\;=\;\underset{y}{\arg\max}\;\sum_{\mathcal{D}_{w}}\mathbf{1}\bigl[y\in y_{w}^{tr}(\bm{x}_{w},y_{w})\bigr],\quad(\bm{x}_{w},y_{w})\in\mathcal{D}_{w},$ (7)
@@ IV-E Efficient Watermark Mitigation
In this stage, CEAT2I combines the benign and watermarked samples identified in the first stage with the triggers extracted in the second stage, and applies a closed-form concept erasure to the fine-tuned model. This process effectively suppresses the watermark while preserving the overall model performance.
@@
Algorithm 1 The pipeline of CEAT2I
Input: A watermarked dataset $\mathcal{D}_{wm}$, the identified watermarked samples $\mathcal{D}_{w}$, the identified benign samples $\mathcal{D}_{b}$, a conditional denoising network $\epsilon_{\bm{\theta}}$, a text encoder $\mathcal{T}$, an image encoder $\mathcal{E}$, an image decoder $\mathcal{R}$, the number of layers in a T2I model $N$, the length of tokens in one sample $L$.
1:Stage 1: Watermarked Sample Detection
2:Obtain a fine-tuned T2I model at an early epoch $T_{e}$
3:for $(\bm{x},y)$ in $\mathcal{D}_{wm}$ do
4:  $\bm{z}=\mathcal{E}(\bm{x}),\bm{c}=\mathcal{T}(y)$
5:  for $i$ = 11 to NN do
6:   Calculate feature deviation $\mathcal{L}_{f}^{i}$ at layer $i$ (Eq. 3)
7:   if $\sum_{i=1}^{N}\mathbf{1}\{\mathcal{L}_{f}^{i}>\alpha_{1}\}>\alpha_{2}$ then
8:     $(\bm{x},y)\rightarrow(\bm{x}_{w},y_{w})\in\mathcal{D}_{w}$
9:   else
10:     $(\bm{x},y)\rightarrow(\bm{x}_{b},y_{b})\in\mathcal{D}_{b}$
11:   end if
12:  end for
13:end for
14:Stage 2: Trigger Identification
15:Obtain a fine-tuned T2I model at a total epoch of $T_{total}$
16:for $(\bm{x}_{w},y_{w})$ in $\mathcal{D}_{w}$ do
17:  Tokenize $y_{w}=\{y_{w}^{1},y_{w}^{2},\dots,y_{w}^{L}\}$
18:  for $i$ = 11 to $L$ do
19:   Calculate feature deviation $\mathcal{L}_{tr}^{i}$ for each token-removal variant $y_{w}\setminus y_{w}^{i}$ (Eq. 5)
20:  end for
21:  Obtain the outlier tokens as the trigger $y_{w}^{tr}$ (Eq. 6)
22:end for
23:Obtain the most frequently occurring outlier tokens as the final trigger $\hat{y}_{w}^{tr}$ for each watermarked sample (Eq. 7)
24:Stage 3: Efficient Watermark Mitigation
25:Obtain a fine-tuned T2I model at a total epoch of $T_{total}$
26:Edit the model using $\mathcal{D}_{w}$, $\mathcal{D}_{b}$, $\hat{y}_{w}^{tr}$ (Eq. 10)
27:return a watermark-free T2I model
@@
By updating the cross-attention weights using this expression, we can effectively erase the model’s sensitivity to specific triggers without degrading its performance on normal inputs. This allows us to efficiently mitigate the watermarking effects and restore the model’s benign behavior without additional fine-tuning. In summary, the algorithm pipeline of our proposed CEAT2I is shown in Algorithm 1.
TABLE I: The CLIP similarity between images (CLIP %) and watermark success rate (WSR %) of one baseline without attacks and five different CEA methods against four types of DOV methods across three datasets, including Pokemon, Ossaili, Pranked03 datasets. The best results among five CEA methods are highlighted in bold. In particular, we mark the failure cases (i.e., WSR $10\%$) among five CEA methods in red.
@@
V Experiments
V-A Main Settings
Datasets and Models. We adopt three benchmark datasets to evaluate all dataset copyright evasion attacks, i.e., Pokemon [45], Ossaili [22], and Pranked03 [23] datasets. For each dataset, we partition the prompts into disjoint training and test sets, using 20% of the data as the test set and the remaining 80% as the training set. All experiments are conducted using Stable Diffusion v1.4, as our default T2I model.
Settings for DOV. We conduct four backdoor-based dataset ownership verifications, including BadT2I-Local (BadT2I-L) [70], BadT2I-Global (BadT2I-G) [70], Rickrolling [55], and Villan Diffusion (VD) [6]. Notably, the threat models underlying these backdoor attacks assume that attackers can fully control the fine-tuning process and thus leverage auxiliary regularization terms to enhance attack performance. In contrast, our study adheres to the DOV setting, in which the data owner is limited to providing the dataset and cannot modify the fine-tuning objective. As a result, these methods cannot be directly applied to DOV, and we adapt them accordingly for our evaluation. Specifically, in all DOV experiments, diffusion models are fine-tuned without any auxiliary regularization terms, ensuring that the observed watermarking behavior stems solely from dataset-level manipulation. Specifically, for the text trigger, BadT2I-L and BadT2I-G use the word “university” as the trigger. Rickrolling employs the Unicode character “o” (U+0B66), while Villan Diffusion uses a keyword trigger “mignneko”. For the owner-specified target image, BadT2I-L is a $128\times 128$ local patch placed at the top-left corner of generated images. BadT2I-G and Rickrolling use a $512\times 512$ global target image, i.e., a Hello Kitty image, while VD uses a $512\times 512$ global target image, i.e., a BabyKitty image. The watermarking rate is set as $\gamma=20\%$. We fully fine-tune the T2I diffusion models on these datasets by using Adam optimizer with a learning rate of $10^{-6}$ for $T_{total}=100$ epochs. The resolution of the generated image is $512\times 512$.
Settings for CEA. We compared our CEAT2I with four different dataset copyright evasion attacks, including ABL [32], NAD [33], TPD [5], and T2IShield [61]. ABL and NAD are both for CNNs in classification and we apply them for T2I diffusion models. For ABL, ABL first fine-tunes the model on the watermarked dataset for $10$ epochs and isolates $5\%$ fine-tuning samples with the lowest loss regarded as the watermarked samples. Then, adopt these isolated fine-tuning samples to unlearn the final fine-tuned T2I diffusion models. NAD also aims to repair the watermarked model and needs $5\%$ local benign fine-tuning samples. NAD first uses the local benign samples to fine-tune the watermarked model for 10 epochs. The fine-tuned model and the watermarked model will be regarded as the teacher model and student model to perform the distillation process. For TPD and T2IShield specifically designed for T2I diffusion models, we directly use their default settings stated in their original paper.
Our CEAT2I performs watermarked sample detection at the early fine-tuning epoch $T_{e}=30$ and the detection thresholds are set to $\alpha_{1}=0.4$ and $\alpha_{2}=15$. We use layers to compute feature differences in the first stage of watermarked sample detection and the second stage of trigger identification. In the first stage, the watermarked sample detection uses all layers of the U-net to calculate feature deviations following [3]. The specific layers used are listed in Appendix. In the second stage, the trigger identification is conducted using the layer that exhibits the largest difference in the average feature deviation $\mathcal{L}_{f}^{i}$ between watermarked and benign samples (i.e., the second-to-last convolutional layer of the model).
Evaluation Metrics. To evaluate the effectiveness of our dataset ownership evasion attacks, we adopt two key metrics from [70]. Specifically, we train a ResNet18 classifier from [70] for each owner-specified target image to detect whether a generated image contains the backdoor-based watermark. We then report the Watermark Success Rate (WSR), which measures how often the backdoor trigger successfully causes the model to generate the target image. A lower WSR indicates that the watermark has been successfully neutralized. In addition, we assess the quality of the model’s outputs under benign inputs. To this end, we compute the CLIP similarity score as [70], which is the cosine similarity between the CLIP embeddings of the generated images and their corresponding ground-truth images. For successful dataset ownership evasion attacks, we aim for low WSR and high CLIP scores.
@@ V-B Main Results
V-C Ablation Study
Results on Watermarked Sample Detection. We compare the effectiveness of watermarked sample detection across ABL [32], T2IShield [61], and our proposed CEAT2I. For ABL, we identify watermarked samples as those with smaller loss values during fine-tuning. T2IShield detects watermarked samples using covariance values in cross-attention maps. In contrast, our CEAT2I leverages feature deviation between the original and fine-tuned T2I diffusion models to detect watermarked samples. Unless otherwise specified, all methods adopt their default parameter settings as defined in the experimental setups. As shown in Table II, ABL achieves low detection accuracy. The limitation arises because T2I diffusion models exhibit highly smooth loss landscapes [66, 18], which weaken the discriminative power of loss-based separation between watermarked and benign samples. T2IShield struggles to detect watermarked samples in BadT2I-L, where the owner-specified target is a small image patch. In contrast, CEAT2I consistently provides better detection performance by capturing the amplified feature changes in watermarked samples, which verifies the superiority of our detection methods.
Results on Trigger Identification. We evaluate the accuracy of our trigger identification approach. Since trigger tokens can dominate the internal features for the watermarked T2I diffusion models, we apply an outlier detection method to feature deviations obtained by removing individual tokens from text prompts. Our CEAT2I method successfully identifies trigger tokens for four DOV methods across three datasets, achieving 100% accuracy when applied to previously detected watermarked samples.
Figure 8: The difference of the average feature deviation $\mathcal{L}_{f}^{i}$ between watermarked and benign samples for each layer against four DOV methods on the Pokemon dataset.
TABLE III: The CLIP similarity between images (CLIP %) and watermark success rate (WSR %) of one baseline without attacks and five different CEA methods against four types of DOV methods on the Pokemon dataset under different watermarking rates. The best results among five CEA methods are highlighted in bold. In particular, we mark the failure cases (i.e., WSR $10\%$) among five CEA methods in red.
@@ V-C Ablation Study
Ablation on the Chosen Layer. In the second stage, the trigger identification uses the layer that exhibits the largest difference in the average feature deviation $\mathcal{L}_{f}^{i}$ between watermarked and benign samples. We compute token-wise removal deviations at the second-to-last convolutional layer and identify tokens whose removal causes significant deviations as candidate triggers. The rationale for this choice is supported by our empirical observations. As shown in Fig. 8, the adopted layer (i.e., the second-to-last convolutional layer) exhibits the largest difference of the average feature deviation $\mathcal{L}_{f}^{i}$ between watermarked and benign samples across four DOV methods on the Pokemon dataset. This empirical evidence motivates our layer selection for reliable trigger identification.
TABLE IV: The CLIP similarity between images (CLIP %) and watermark success rate (WSR %) of one baseline without attacks and five different CEA methods against four types of DOV methods on the Pokemon dataset under different trigger positions. The best results among five CEA methods are highlighted in bold. In particular, we mark the failure cases (i.e., WSR $10\%$) among five CEA methods in red.
@@
TABLE V: The CLIP similarity between images (CLIP %) and watermark success rate (WSR %) of one baseline without attacks and five different CEA methods on the Pokemon dataset under a single word trigger and a phrase trigger. The best results among five CEA methods are highlighted in bold. In particular, we mark the failure cases (i.e., WSR $10\%$) among five CEA methods in red.
Trigger Target Image No Attack ABL NAD TPD T2IShield CEAT2I (Ours)
CLIP WSR CLIP WSR CLIP WSR CLIP WSR CLIP WSR CLIP WSR
university Local Patch 89.5 85.6 78.2 81.8 85.7 17.8 89.5 90.8 81.2 25.8 89.5 4.6
university Global HelloKitty 89.7 84.5 80.7 85.4 88.1 53.1 90.9 72.8 81.5 7.6 90.0 3.2
o (U+0B66) Global HelloKitty 90.1 99.7 73.1 94.2 78.2 50.9 85.3 60.1 84.3 12.2 89.7 1.6
university Global BabyKitty 89.7 99.7 78.2 95.3 88.0 63.2 89.2 70.2 85.2 10.8 89.5 2.5
dataset copyright protection Local Patch 89.3 88.5 76.2 86.8 80.1 80.4 87.1 87.4 85.5 58.3 90.2 3.9
Global HelloKitty 88.8 99.9 80.0 99.7 87.9 97.8 89.2 99.3 84.7 8.6 89.5 3.5
Global BabyKitty 89.7 99.8 90.6 96.9 84.7 98.5 85.7 99.7 88.8 13.9 89.7 2.9
TABLE VI: The computational time (h) of one baseline without attacks and five different CEA methods against BadT2I-L on Pokemon, Ossaili, and Pranked03 datasets. For the five CEA methods, the reported time cost reflects only the additional computational overhead beyond the fine-tuning stage, since model fine-tuning on watermarked datasets is required even under the no-attack setting.
Dataset No Attack ABL NAD TPD T2IShield CEAT2I (Ours)
Stage 1 Stage 2 Stage 3 Total
Pokemon 8.5 1.6 1.9 0.0 0.9 0.5 0.4 0.3 1.2
Ossaili 12.0 2.2 2.4 0.0 1.2 0.6 0.5 0.3 1.4
Pranked03 60.0 11.2 13.5 0.0 6.1 3.0 2.8 1.5 7.3
@@
V-D Discussions
Discussions on Single-Word Trigger and Phrase Trigger. We explore the impact of both single-word and phrase triggers using the Pokemon dataset. Specifically, BadT2I-L, BadT2I-G, and VD use the single-word trigger “university” with different target images, while Rickrolling uses the single-character trigger “o (U+0B66)” in default. Additionally, we adopt the phrase “dataset copyright protection” as a multi-word trigger. As shown in Table V, CEAT2I remains highly effective for phrase triggers and achieves the lowest WSRs compared to baselines. This is primarily because trigger words dominate the learned features, and even when the trigger appears as a phrase, each constituent word behaves as a statistical outlier. Consequently, the proposed outlier-based detection in the second stage can effectively capture such phrase triggers.
Discussions on Computational Cost. In Table VI, we report the computational time (in hours) of one baseline without attacks and five different CEA methods against BadT2I-L across three datasets. All the experiments are conducted on one NVIDIA 4090 GPU. For the five CEA methods, the reported time cost represents only the additional computational overhead beyond the mandatory fine-tuning stage, since fine-tuning on watermarked datasets is required even in the no-attack setting. As shown in this table, among these methods, TPD incurs almost no additional time cost, as it only applies random perturbations to input texts during inference. In addition, CEAT2I requires similar time as T2IShield but less than ABL and NAD, as both ABL and NAD involve additional fine-tuning of the watermarked models. Notably, CEAT2I achieves the lowest WSRs among all compared approaches, demonstrating its superior effectiveness. In the future work, we plan to investigate strategies for further reducing the computational overhead of CEAs while preserving its effectiveness.
Discussions on Watermarks for Style Protection. In addition to the backdoor-based DOV, we evaluate the robustness of a representative watermarking approach for style protection, SIREN [30], under our CEAT2I. SIREN embeds an imperceptible but learnable coating into protected fine-tuning datasets so that personalized diffusion models can reliably capture it during the fine-tuning process. For verification under black-box conditions, features of generated outputs are extracted and classified to determine the presence of the coated signature.
Following SIREN, we conduct experiments on five datasets, including Pokemon [45], CelebA-HQ [26], ArtBench [38], Landscape [25], and WikiArt [53] datasets. Unless otherwise specified, the fine-tuning and evaluation configurations follow the original SIREN paper. For robustness evaluation, we use a baseline without attacks and our CEAT2I described in our original manuscript under the same settings. Effectiveness is evaluated using Bit Accuracy (BitAcc) for verification and CLIP scores for generation quality. The higher BitAcc and CLIP scores indicate better reliability of the method.
We compare the BitAcc and CLIP scores of a baseline without attacks and our CEAT2I against SIREN across five datasets, as reported in Table VII. In the absence of attacks, SIREN attains high BitAcc for verification while preserving strong generation quality on benign images. However, SIREN suffers significant performance degradation under our CEAT2I, as evidenced by notably reduced BitAcc and CLIP scores. These results highlight that our proposed CEAT2I effectively targets both harmful backdoor-based watermarking methods (BadT2I-G, Rickrolling, and VD) and harmless DOV schemes (BadT2I-L and SIREN).
V-E Resistance to Potential Adaptive Defense
@@
TABLE VII: The CLIP similarity between images (CLIP %) and Bit Accuracy (BitAcc %) of one baseline without attacks and our CEAT2I against SIREN on five datasets, including Pokemon, CelebA-HQ, ArtBench, Landscape, WikiArt.
Dataset No Attack CEAT2I (Ours)
CLIP BitAcc CLIP BitAcc
Pokemon 88.2 94.5 86.4 0.0
CelebA-HQ 86.8 95.0 88.5 0.0
ArtBench 87.5 92.4 85.0 0.0
Landscape 89.0 96.5 86.4 0.0
WikiArt 85.5 90.0 83.6 0.0
Let $f_{\bm{\theta}}^{i}\left(\bm{z}_{t},t,\bm{c}\right)$ and $f_{\bm{\theta}_{\bm{w}_{t}}^{\prime}}^{i}\left(\bm{z}_{t},t,\bm{c}\right)$ denote the feature activations at the $i$-th layer of the original and fine-tuned diffusion models at an early fine-tuning epoch $T_{e}$, respectively. For an image $\bm{x}$ and text input $y$, we define $\bm{z}_{t}=\mathcal{E}(\bm{x})$ as the latent representation at diffusion timestep $t$, and $\bm{c}=\mathcal{T}(y)$ as the semantic text embedding. Let $N$ be the total number of layers in the diffusion model. We denote benign image–text pairs by $(\bm{x}_{b},y_{b})\in\mathcal{D}_{b}$, and watermarked pairs by $(\bm{x}_{w},y_{w})\in\mathcal{D}_{w}$, where $\bm{x}_{w}$ is the target image and the watermarked prompt is $y_{w}=y\oplus p$, with $p$ representing the trigger tokens. In the adaptive attack, the attacker optimizes a trigger $p$ such that the watermarked prompt $y_{w}$ produces a target image $\bm{x}_{w}$ while minimizing the feature deviation of the watermarked and benign samples. This objective can be formalized as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min_{p}\frac{1}{N|\mathcal{D}_{w}|}\sum_{(\bm{x}_{w},y_{w})\in\mathcal{D}_{w}}\sum_{i=1}^{N}||f_{\bm{\theta}}^{i}\left(\mathcal{E}(\bm{x}_{w}),t,\mathcal{T}(y_{w})\right)}$ (11)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-f_{\bm{\theta}_{\bm{w}_{t}}^{\prime}}^{i}\left(\mathcal{E}(\bm{x}_{w}),t,\mathcal{T}(y_{w})\right)||_{2}^{2}}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{1}{N|\mathcal{D}_{b}|}\sum_{(\bm{x}_{b},y_{b})\in\mathcal{D}_{b}}\sum_{i=1}^{N}||f_{\bm{\theta}}^{i}\left(\mathcal{E}(\bm{x}_{b}),t,\mathcal{T}(y_{b})\right)}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-f_{\bm{\theta}_{\bm{w}_{t}}^{\prime}}^{i}\left(\mathcal{E}(\bm{x}_{b}),t,\mathcal{T}(y_{b})\right)||_{2}^{2}.}$